Draw all possible random samples of size 2 with replacement from the
population 2, 3, 4, 5 and 7. Similarly draw all possible random samples of size
2 with replacement from other population 1, 2, 2, 4 and 5.Find the sampling
distribution of difference b/w two sample means and calculate its Mean,
Variance and Standard Error. Also find the Mean, Variance and Standard
Deviation of two Populations and Verify the Results.
1
Expert's answer
2020-09-28T21:12:35-0400

solution


when drawn with replacement, the total number of combinations that can be obtained is:


"n^r"

Where r is the number of elements chosen and n the number of elements to Choose from.


"=5^2=25"



From the population "X= (2, 3, 4, 5, 7)" each of the samples composed of 2 Numbers has a mean of:


"\bar x_i = \frac{a_i+b_i}{2}"

Where "\bar x_i" Is the mean of the "i^{th}" Combination. A series of sample means is formed i.e.

"X_{means} = \bar x_1, \bar x_2, \bar x_3,...,\bar x_{25}"


From the population"Y=(1,2,2,4,5)" each of the samples composed of 2 Numbers has a mean of:


"\bar y_i = \frac{a_i+b_i}{2}"

Where "\bar y_i" Is the mean of the "i^{th}" Samples. A series of sample means is formed i.e


"Y_{means} = \bar y_1, \bar y_2, \bar y_3,...,\bar y_{25}"

The distribution of differences between the 2 means is given by:



"differences =X_{means}-Y_{means}"

The distribution of mean differences is obtained as:


"(difference,\ frequency)=\ (1,9),\ (1.5,12),\ (2,4)"

1. Answers for the distribution of differences in means


Mean:


"\bar {difference} =\frac{ \sum (differences * frequency)}{n}""=\frac{(1*9)+(1.5*12)+(2*4)}{25} = 1.4"

mean =1.4

Variance:


"\sigma^2 =\frac{ \sum frequency *
 (differences -\bar {difference})}{n-1} = 0.29167"

variance = 0.125


standard error:


"SE = \sqrt {\sigma^2} = \sqrt{0.125} = 0.3536"

standard error = 0.5401


2. answers for the distribution of X and Y


Mean of the population X:


"\frac{\sum X}{n}=\frac{2+3+4+5+7}{5}=4.2"

mean of X = 4.2


Variance of the distribution of X:



"\sigma ^2 = \frac{\sum( X - \bar X)^2
}{n-1} = 3.7"

variance of X = 3.7


Standard error of the distribution of means of X:


"\sqrt {\sigma^2} = \sqrt{3.7} = 1.924"

standard deviation of X= 1.924


The mean of the population Y:


"\frac{\sum Y}{n}=\frac{1+2+2+4+5}{5}=2.8"



mean of Y = 2.8


Variance of the distribution of means of Y:


"\sigma ^2 = \frac{\sum( Y- \bar Y)^2
}{n-1} =2.7"

variance of Y = 2.7


Standard error of the distribution of means of Y:


"\sqrt{\sigma^2}=\sqrt{2.7}= 1.6432"


standard deviation of Y = 1.6432


The difference between the mean on X and mean of Y:


"\bar X- \bar Y = 4.2-2.8 = 1.4" . This is equal to the mean of the distribution of differences in means ("\bar {difference}" )